# Non-Equilibrium Aspects of Fission Dynamics within the Time Dependent Density Functional Theory

### v1

@@ Abstract
In this report we will cover in some details only our preliminary results for the induced fission of odd-mass and odd-odd nuclei within a time-dependent density functional (TDDFT) extended to superfluid systems. In the talk we have covered three topics; i) Scission neutrons, ii) Induced fission of odd-mass and odd-odd nuclei, and iii) Nontrivial aspects of fission dynamics, which we discuss here briefly.
@@ 1 Introduction
In another relatively recent publication Bulgac:2024a we have presented several new aspects of the fission process, which were never discussed in literature and this topic will not be presented here in more detail. In particular we have presented evidence that the fission dynamics from saddle-to-scission has an intrinsically non-Markovian character and consequently various previous attempts to describe this process using various stochastic or kinetic approaches are inadequate. Moreover, since in nuclear large amplitude collective dynamics (LACM) the occupation probabilities of single-particle states with large momenta have very long tails $\propto C/p^{4}$, where $p$ is the local single-particle momentum, a significant repopulation of the quasiparticle states occurs during the fission dynamics and theoretical modeling with a reduced basis set of single-particle states, such as (TD)BCS+TDHF or even TDDFT with a limited number of quasiparticle states lead to inaccurate predictions of the fission fragments (FFs), or even fail to lead to fission.
The next section will cover only a broad overview of our new results concerning the induced fission of odd-mass and odd-odd nuclei, which have not been submitted yet for publication and are still under analysis.
Figure 1: In the upper row left panel we display the time difference to full spatial separation of the two FFs between the case of an even-even and odd-mass nuclei. The odd-mass system is treated with an exact odd-neutron and even-proton numbers respectively. The odd-mass nucleus is treated as having an even-neutron and an even-proton numbers, with the average $N=147$ and $Z=92$ respectively for 239U and then a quasiparticle state is excited as described in Refs. Ring:2004 ; Bertsch:2009 . In the upper row right panel we show several TDDFT trajectories for the fission of an odd-nucleon nucleus with solid lines and for reference a TDDFT trajectory for an even-even nucleus with average $N=147$ and $Z=92$, with a dashed-thin solid line. Unlike the TDDFT trajectories, in which typically the quadrupole moment $Q_{20}$ typically increases in time in the case of fission of odd-nucleon systems the nucleus encounters a “bumpier” potential energy surface in the $(Q_{20},Q_{30})$ plane towards scission and the system it gets “confused” following a much more "challenging obstacle course," see upper row right panel. In the second row we show the occupation probability of the initial odd-nucleon state as a function of time. For time $t=0$ the occupation probability was obtained by constructing the corresponding canonical occupation probabilities Bulgac:2023 . For $t>0$ apart from one canonical quasiparticle state with occupation probability 1, the rest of the quasiparticle states are all double occupied with probabilities less than 1.
2 Induced fission of odd-mass and odd-odd nuclei
Ever since the development of a microscopic description of pairing correlations by Bardeen et al. Bardeen:1957 and the recognition that they are also most likely crucial in explaining nuclear properties Bohr:1958 , the treatments of fermion systems with odd numbers of particles have relied on the so called Pauli Blocking approximation. The extra odd nucleon carries a non-vanishing angular momentum and as a result the nuclear mean field brakes time-reversal invariance and the entire nucleus as well has a negative time-parity. From the theoretical point of view this is a rather inconvenient feature, which brings in quite a number of technical difficulties in the treatment of nuclear fission, and which forced most theoretical treatments to adopt a number of approximations and assumptions. In nuclear LACM the Pauli blocking approximation has been used almost always, see the most recent and detailed studies of the fission of odd-mass nuclei, where Pauli blocking in conjunction along with the equal filling approximation Perez-Martin:2008 have been used Schunck:2023 ; Pore:2024 .
We will present briefly here an overview of our results in the study of the fission of odd nuclei 239U (23 trajectories), 241Pu (42 trajectories), 243Pu (13 trajectories), and the odd-odd nucleus 238Np (35 trajectories), where in each case the number of trajectories describe various initial nuclear shape deformations characterized by quadruple and octupole momenta $(Q_{0},Q_{30})$ and various odd quasiparticle states of the respective fissioning nucleus, constructed using the prescription outlined in Refs. Ring:2004 ; Bertsch:2009 ; Schunck:2023 ; Pore:2024 . In total we have accumulated 113 trajectories for these nuclei, all obtained within the TDDFT extended to superfluid systems, a number comparable (or maybe even bigger) with perhaps of all of the rest of the world TDDFT fission trajectories obtained for even-even nuclei by other authors since 2015. The simulations have been performed in a simulation box $30^{2}\times 60$ and following the dynamics for up to about 60,000 times steps or longer in some cases. These simulations were performed with the energy density functional SeaLL1 Bulgac:2018 by evolving $0002\times 2\times 4\times 30^{2}\times 60=864,$ 000 complex, coupled, nonlinear PDEs in 3D+time, see Refs Bulgac:2019c ; Bulgac:2020 ; Bulgac:2024a for details.
Overall the FF properties in the induced fission of even-even, odd-mass, and odd-odd nuclei are very similar. Often there is a difference in the times the two FFs reach scission, when the corresponding initial quasiparticle state has a relatively low excitation energy, typically smaller than 1 MeV, see Fig. 1. In that case the TDDFT trajectory is significantly more complicated as the system spends quite a lot of time ’deciding’ which path to follow on the potential energy surface. Such a situation has been encountered for some almost symmetric TDDFT trajectories Bulgac:2019c ; Bulgac:2020 in the case of fission even-even compound nuclei.
The most unexpected feature is however the violation of the Pauli blocking approximation, see Fig. 1, used for more than 6 decades in nuclear physics. During the time evolution, as was demonstrated in Ref. Bulgac:2024a , even if the system is started in a fully canonical basis, in the subsequent time evolution the quasiparticle states are not anymore canonical states, and at each time step one has to identify the new canonical set of wave functions. Only in this case one can clearly see that the single odd quasiparticle state occupied with probability 1, ceases to maintain the same occupation probability. However, at each time $t>00$ there is only one canonical state occupied with exactly probability 1. We expect that this is a feature to be observed in any nuclear LACM of an odd fermion system, in either excited states of a nucleus or in heavy-ion collisions. A clear example when such a situation is encountered is when the angular momentum of the initial odd fermion is greater than $\approx A^{1/3}$, since such a state cannot exist with a significant occupation probability in either FF Bertsch:1997 ; Bulgac:2016 ; Bulgac:2019c ; Bulgac:2020 .
@@ 2 Induced fission of odd-mass and odd-odd nuclei
The funding for A.B. and M.K. from the Office of Science, Grant No. DE-FG02-97ER41014 and also the partial support provided by NNSA cooperative Agreement DE-NA0003841 is greatly appreciated. The work of I.A. and I.S. was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. The Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy Contract No. 89233218CNA000001. This research used resources of the Oak Ridge Leadership Computing Facility, which is a U.S. DOE Office of Science User Facility supported under Contract No. DE-AC05-00OR22725. The report is published as a preprint with numbers NT@UW-24-11 and LA-UR-24-xxxxxx.

### v2

@@ Abstract
We will cover briefly the time-dependent density functional theory approach, extended to include pairing correlations, to induced fission of 238U(n,f), 241,243Pu, and 238Np, and also results on scission neutrons and a number of very nontrivial aspects of induced fission.
@@ 1 Introduction
For all results reported here we performed extensive fully microscopic studies of fission dynamics within the time-dependent density functional theory (TDDFT) extended to include pairing correlations, with controlled numerical approximations and without introducing any unchecked assumptions or simplifications. The theoretical framework is described and was reviewed in great detail in Refs. Bulgac:2019 ; Shi:2020 ; Bulgac:2018 and the corresponding codes are accessible on GitHub.
2 Induced fission of odd-mass and odd-odd nuclei
We performed studies of the induced fission dynamics of nuclei with either odd number of neutrons for the compound fission systems 238U(n,f) [23 trajectories], 240Pu(n,f) [42 trajectories], 242Pu(n,f) [7 trajectories], and with both odd proton and neutron numbers for 237Np(n,f) [48 trajectories], where in each case the number of trajectories describe various initial nuclear shape deformations characterized by quadrupole and octupole momenta $(Q_{20},Q_{30})$ near the outer fission barrier and various odd quasiparticle states of the respective fissioning compound nucleus, constructed using the standard prescription outlined in Refs. Ring:2004 ; Bertsch:2009 ; Schunck:2023 ; Pore:2024 . In addition, for comparison, we have performed additional calculations also for the induced fission of 235U(n,f) [24 trajectories], 239Pu(n,f) [7 trajectories] and for the odd nucleon systems 237Np(n,f) [11 trajectories], 238U(n,f) [6 trajectories], 240Pu(n,f) [13 trajectories], and 242Pu(n,f) [2 trajectories], treating them as even-even nuclei for a total of 63 trajectories. The total 121+63 = 184 TDDFT trajectories for these nuclei is comparable (or maybe even bigger) with perhaps of all of the rest of the world TDDFT fission trajectories obtained for even-even nuclei by other authors since 2015, see review talk of D. Vretenar for some references. The simulations have been performed in a simulation box $30^{2}\times 60$ fm3 with a lattice constant of 1 fm and following the dynamics for up to about 30,000 to 60,000 times steps or longer in some cases. These simulations were performed with the energy density functional SeaLL1 Bulgac:2018 by evolving $0002\times 2\times 4\times 30^{2}\times 60=864,$ 000 complex, coupled, nonlinear PDEs in 3D+time, see Refs Bulgac:2018 ; Bulgac:2019 ; Shi:2020 ; Bulgac:2019c ; Bulgac:2020 ; Bulgac:2023 ; Bulgac:2024a for details. This proves that inspite of the complexity of these simulations, using a total number of quasiparticle states $0002\times 2\times 30^{2}\times 60=216,$ 000, such simulations can be performed on a large variety of supercomputers available worldwide nowadays. Detailed extensive results of these simulations will be submitted for publication soon.
Figure 1: In the upper row left panel we display the time difference to full spatial separation of the two fission fragments (FFs) between the case of an even-even and odd-mass nuclei. The odd-mass system is treated with an exact odd-neutron and even-proton numbers respectively. The odd-mass nucleus is treated as having an even-neutron and an even-proton numbers, with the average $N=147$ and $Z=92$ respectively for 239U and then a quasiparticle state is excited as described in Refs. Ring:2004 ; Bertsch:2009 . In the upper row right panel we show several TDDFT trajectories for the fission of an odd-nucleon nucleus with solid lines and for reference a TDDFT trajectory for an even-even nucleus with average $N=147$ and $Z=92$, with a dashed-thin solid line. Unlike the TDDFT trajectories, in which typically the quadrupole moment $Q_{20}$ typically increases in time in the case of fission of odd-nucleon systems the nucleus encounters a “bumpier” potential energy surface in the $(Q_{20},Q_{30})$ plane towards scission and the system it gets “confused” following a much more "challenging obstacle course," see upper row right panel. In the second row we show the occupation probability of the initial odd-nucleon state as a function of time. For time $t=0$ the occupation probability was obtained by constructing the corresponding canonical occupation probabilities Bulgac:2023 . For $t>0$ apart from one canonical quasiparticle state with occupation probability 1, the rest of the quasiparticle states are all double occupied with probabilities less than 1.
Ever since the development of a microscopic description of pairing correlations by Bardeen et al. Bardeen:1957 and the recognition that they are also most likely crucial in explaining nuclear properties Bohr:1958 , the treatments of fermion systems with odd numbers of particles have relied on the so called Pauli blocking approximation. The extra odd nucleon carries a non-vanishing angular momentum and as a result the nuclear mean field breaks time-reversal invariance. The compound nucleus breaks the gauge symmetry as well and has a negative particle parity. The rotational symmetry and sometimes spatial parity are also broken. From the theoretical point of view these are rather inconvenient features, resulting in a significant number of technical difficulties, and which forced most theoretical treatments to adopt approximations and assumptions. In nuclear large amplitude collective motion the Pauli blocking approximation has been used almost always, see the most recent and detailed studies of the fission of odd-mass nuclei, where Pauli blocking in conjunction along with the equal filling approximation Perez-Martin:2008 have been implemented Schunck:2023 ; Pore:2024 .
Overall the FF properties in the induced fission of even-even, odd-mass, and odd-odd nuclei are very similar, but the widths of total kinetic energy, FF neutron and charge distributions are typically twice as large as in the case of induced fission of even-even nuclei. Often there is a difference in the times the two FFs reach scission, when the corresponding initial quasiparticle state has a relatively low excitation energy, typically smaller than the pairing gap, see Fig. 1. In that case the TDDFT trajectory is significantly more complicated as the system spends quite a lot of time “deciding”’ which path to follow on the potential energy surface. Such a situation has been encountered for some almost symmetric TDDFT trajectories Bulgac:2019c ; Bulgac:2020 in the case of fission even-even compound nuclei.
An unexpected feature is the violation of the Pauli blocking approximation, see Fig. 1, used for more than 6 decades in nuclear physics. During the time evolution, as was demonstrated in Ref. Bulgac:2024a , even if the system is started in a fully canonical basis, in the subsequent time evolution the quasiparticle states are not anymore canonical states, and at each time step one has to identify the new canonical set of wave functions. One can clearly see that the single odd quasiparticle state initially occupied with probability 1 in canonical basis, ceases to maintain the same occupation probability. However, at each time $t>00$ there is only one canonical state occupied with exactly probability 1. We expect that this is a feature to be observed in any nuclear Large Amplitude Collective Motion (LACM) of an odd fermion system, in either excited states of a nucleus or in heavy-ion collisions. A clear example when such a situation is encountered is when the angular momentum of the initial odd fermion is $\approx A^{1/3}$, since such a state cannot exist with a significant occupation probability in either FF Bertsch:1997 ; Bulgac:2016 ; Bulgac:2019c ; Bulgac:2020 .
3 Scission neutrons and nontrivial aspects of fission dynamics revealed in TDDFT framework
@@
In recent publications Bulgac:2023 ; Bulgac:2024a we have presented several new aspects of the fission process, which were never discussed in literature. We have demonstrated that the fission dynamics from saddle-to-scission has an intrinsically non-Markovian character and consequently various previous attempts to describe this process using various stochastic or kinetic approaches are inadequate. Moreover, since in nuclear LACM the occupation probabilities of single-particle states with large momenta have very long tails $\propto C/p^{4}$, where $p$ is the local single-particle momentum, a significant repopulation of the quasiparticle states occurs during the fission dynamics and theoretical modeling with a reduced basis set of single-particle states, such as (TD)BCS+TDHF or even TDDFT with a limited number of quasiparticle states lead to inaccurate predictions of the FFs properties, or even fail to lead to fission.
4 Conclusions
TDDFT, extended to include a correct description of the dynamics of the pairing correlations, is arguably the most advanced microscopic quantum tool available today, and it is based on solid independent knowledge of basic nuclear properties of nuclei with $A\geq 16$. In this approach only 7 parameters well known for a long time are needed: energy and equilibrium density of symmetric nuclear matter, surface tension, symmetry energy and its density dependence, spin-orbit and pairing strengths), and this leads to a one of the most accurate descriptions of the binding energies, charge radii, one- and two-neutron separation energies, single-particle spectra of magic nuclei, etc., see Ref. Bulgac:2018 , and it is used in TDDFT to describe fission dynamics. We described briefly several recent developments in a fully microscopic description of fission dynamics. i) The TDDFT approach to the emission of scission neutrons, a topic still not settled either theoretically or experimentally, was conjectured in 1939 Bohr:1939 soon after nuclear fission explanation of fission Meitner:1939 . ii) The non-Markovian aspects of the fission descent, which is distinct from the earlier demonstrated clear dissipative character of fission dynamics. iii) A first peak to some of our preliminary results from the first TDDFT treatment of the induced fission of a large class of odd nucleon compound nuclei. A overall arching conclusion emerges from these studies: fission dynamics is an intrinsically non-equilibrium process, which requires a detailed time-dependent microscopic description, without the inclusion of unchecked assumptions, phenomenology and parameter fitting, and uncontrolled approximations and simplifications.
@@ 4 Conclusions
The funding for A.B. and M.K. from the Office of Science, Grant No. DE-FG02-97ER41014 and also the partial support provided by NNSA cooperative Agreement DE-NA0003841 is greatly appreciated. The work of I.A. and I.S. was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. The Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy Contract No. 89233218CNA000001. This research used resources of the Oak Ridge Leadership Computing Facility, which is a U.S. DOE Office of Science User Facility supported under Contract No. DE-AC05-00OR22725.